# A note on the triple product property for finite groups with abelian normal subgroups of prime index

### v5

@@ 5. Analysis and Conclusion
In the case of $p=2$ and the dihedral groups $D_{2n}$ the theorem proves that $\rho_{0}(D_{2n})\leq\frac{4}{3}$ [5, Conjecture 7.5], and together with the data for TPP triples in groups of order $\leq 32$ obtained via exhaustive computational search [5, Table 1], strengthens the conjecture that $\rho\leq\frac{4}{3}$ for groups with cyclic (and normal) subgroups of index $2$ [5, Conjecture 7.6].
For $p>2$ the smallest known group containing a cyclic normal subgroup of index $p$ and a known $\rho$ is one of type $C_{7}\rtimes C_{3}$ (GAP ID [21, 1]), a non-abelian group of order $21=3\cdot 7$ containing a cyclic normal subgroup $C_{7}$ of index $p=3$, and which realises a largest TPP triple of type $(3,3,3)$ and size $3^{3}=27$, and satisfies the bound $\frac{27}{21}<\frac{3^{2}}{2\cdot 3-1}=\frac{9}{5}=1.8$ [5, Table 1]. The same data also shows that there are actually no groups of order $\leq 32$ containing cyclic normal subgroups of index $3$ and realising $\rho>\frac{9}{5}$. This seems to be true even for larger groups of order up to $100$, as indicated in the results of Xiang et. al. in 2018, who use a computational approach based on evolutionary search algorithms to look at TPP capacity $\beta$ (and other parameters related to group-theoretic matrix multiplication) in a selection of groups of order up to $100$ [8, Tables 1, 3-4]: some simple GAP computations show that of all groups listed in those tables which realise $\rho>\frac{9}{5}$ none contain cyclic normal subgroups of index $3$. Noting that a subgroup of a group which is of smallest prime index is always normal, this suggests the following conjecture.
@@
If $G$ is a group with a cyclic normal subgroup of prime index $p$, and $(S,T,U)$ is any TPP triple of $G$, then $|S||T||U|\leq\frac{p^{2}}{2p-1}|G|$, i.e. $\rho(G)\leq\frac{p^{2}}{2p-1}$. For this bound to be as tight as possible $p$ can be taken to be the smallest such prime, in which case the cyclic subgroup of index $p$ will necessarily be normal.
The conjecture is false for the more general case of abelian normal subgroups of prime index: for $p=2$, where the upper bound to be respected is $\rho\leq\frac{4}{3}$, the smallest known counterexamples are two $2$-groups of order $32$, one of type $(C_{4}\times C_{4})\rtimes C_{2}$ (GAP ID [32, 11]) and another of type $(C_{2}\times C_{2}\times C_{2}\times C_{2})\rtimes C_{2}$ (GAP ID [32, 27]), both containing a non-cyclic abelian subgroup of index $2$, but no cyclic abelian subgroups of index $2$, and realising $\rho=1.5>\frac{4}{3}$ (via TPP triples of type $(6,4,2)$).

### v7

@@ 5. Concluding Remarks
In the case of $p=2$ and the dihedral groups $D_{2n}$ the theorem proves that $\rho_{0}(D_{2n})\leq\frac{4}{3}$ [5, Conjecture 7.5]. The data for TPP triples in groups of order $\leq 32$ obtained via exhaustive computational search [5, Table 1] shows that none of those groups realise $\rho>\frac{4}{3}$ if they contain a cyclic subgroup of index $2$, and this led to the original conjecture that $\rho\leq\frac{4}{3}$ for groups with cyclic subgroups of index $2$ [5, Conjecture 7.6].
The smallest known (nonabelian) group containing a cyclic normal subgroup of index $p>2$ and a known $\rho>1$ is one of type $C_{7}\rtimes C_{3}$ (GAP ID [21, 1]), a group of order $21=3\cdot 7$ containing a cyclic normal subgroup $C_{7}$ of index $p=3$, and that realises a largest TPP triple of type $(3,3,3)$ and size $3^{3}=27$, and satisfies the bound $\frac{27}{21}<\frac{3^{2}}{2\cdot 3-1}=\frac{9}{5}=1.8$ [5, Table 1]. The same data also shows that there are actually no groups of order $\leq 32$ containing cyclic normal subgroups of index $3$ and realising $\rho>\frac{9}{5}$. This seems to be true even for larger groups of order up to $100$, as indicated in the results of Xiang et. al. in 2018, who use a computational approach based on evolutionary search algorithms to look at TPP capacity $\beta$ (and other parameters related to group-theoretic matrix multiplication) in a selection of groups of order up to $100$ [8, Tables 1, 3-4]: some simple GAP computations show that of all groups listed in those tables that realise $\rho>\frac{9}{5}$ none contain cyclic normal subgroups of index $3$. Noting that subgroups of smallest prime index are always normal, this suggests the following conjecture (which is a generalisation of [5, Conjecture 7.6]).
@@
If $G$ is a group with a cyclic normal subgroup of prime index $p$, and $(S,T,U)$ is any TPP triple of $G$, then $|S||T||U|\leq\frac{p^{2}}{2p-1}|G|$, that is, $\rho(G)\leq\frac{p^{2}}{2p-1}$. For this bound to be as tight as possible $p$ can be taken to be the smallest such prime, in which case the cyclic subgroup of index $p$ will necessarily be normal.
The conjecture is false for the more general case of abelian normal subgroups of prime index: for $p=2$, where the upper bound to be respected is $\rho\leq\frac{4}{3}$, the smallest known counterexamples are two $2$-groups of order $32$, one of type $(C_{4}\times C_{4})\rtimes C_{2}$ (GAP ID [32, 11]) and another of type $(C_{2}\times C_{2}\times C_{2}\times C_{2})\rtimes C_{2}$ (GAP ID [32, 27]), both containing a non-cyclic abelian subgroup of index $2$, but no cyclic subgroups of index $2$, and realising $\rho=1.5>\frac{4}{3}$ (via TPP triples of type $(6,4,2)$).